OpenStudy (anonymous):

suppose y varies directly as x. write a direct equation that relates x and y. then solve. if y=45 whenx=15, find x when y=45

OpenStudy (anonymous):

y = k * x 45 = k * 15 , what is k

OpenStudy (anonymous):

k =

OpenStudy (anonymous):

Is the answer 3

OpenStudy (anonymous):

k=3

OpenStudy (anonymous):

find x when y=45

OpenStudy (anonymous):

y = k*x we know: k=3, y=45 45 = 3*x the answer to this problem is what is the value of x

OpenStudy (anonymous):

The answer is y=3x correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks can u help me on another 1

OpenStudy (anonymous):

second part of question: "find x when y=45" we can use our equation: y=3x and set y to 45

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

Thanks can u help me on another 1

OpenStudy (anonymous):

sorry...... you did not solve the second part of the question

OpenStudy (anonymous):

do you know the answer to the second part ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

y=3x if y is 45, we have 45 = 3x

OpenStudy (anonymous):

so x must be 15........ that's answer to second question :)

OpenStudy (anonymous):

"write a direct equation that relates x and y." y=3x "find x when y=45" x=15

OpenStudy (anonymous):

If y=-4 when x=2, find y when x=-6

OpenStudy (anonymous):

we must again make the equation y = k*x

OpenStudy (anonymous):

first, we have x and y -- find k -4 = k * 2

OpenStudy (anonymous):

Is that part -2

OpenStudy (anonymous):

k=-2 :)

OpenStudy (anonymous):

so our equation is: y = (-2) * x we can use the equation to solve second part find y when x=-6

OpenStudy (anonymous):

Is it 3

OpenStudy (anonymous):

looks like it because of the way it is written y=(-2) * (-6) <--- find y

OpenStudy (anonymous):

they had given us x, not y this time

OpenStudy (anonymous):

So is it 12

OpenStudy (anonymous):

yep :) answer is 12

OpenStudy (anonymous):

y is 12, when x=-6 our equation: y = -2 * x

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

If y =-9 when x=3, find y when x=-5

OpenStudy (anonymous):

Would it be y=(-3)×x would it be 15

OpenStudy (anonymous):

yes and yes completely right

OpenStudy (anonymous):

If y=4 when x=16,find y when x=6

OpenStudy (anonymous):

well y is 1/4 of x, so k=1/4 if that is the constant, then y = 1/4 * 6 which is 6/4 or 1 2/4 or 1 1/2

OpenStudy (anonymous):

Okay would it be y=(.25)×x woukd it be 1.5

OpenStudy (anonymous):

yes, exactly

OpenStudy (anonymous):

If y=12 when x=18,find x when y=-16

OpenStudy (anonymous):

now it starts to get difficult y = k*x 12 = k*18

OpenStudy (anonymous):

to get k alone, we must divide both sides by 18: 12/18 = k

OpenStudy (anonymous):

.66

OpenStudy (anonymous):

yeah, k=0.66

OpenStudy (anonymous):

"find x if y=-16": -16 = 0.66 * x

OpenStudy (anonymous):

Would that be -24.24

OpenStudy (anonymous):

both sides multiplied by 100: -1600 = 66 x divide both sides by 66 -1600/66 = x x = -1600/66 = -24.24, yes

OpenStudy (anonymous):

Okay thanks :D ur amazing

OpenStudy (anonymous):

thanks :P

OpenStudy (anonymous):

I wish u could send pictures of the problem cause I dont know how to do it and it is a graph